You need to buy 36,000 pounds of orange juice concentrate in three months. You go short position with the speculator at P54/lbs. If the price of orange juice is P56 after three months.

  1. How much is your gain or loss?
  2. How much is the gain or loss of the speculator?

Price continuously rising so the speculator go short hedge with the hedger at P56. If the price of orange juice is P57 after three months.

  1. How much is the gain or loss of the speculator regarding this transaction?
  2. How much is the gain or loss of the hedger?
  3. How much is the gain or loss of the speculator for the whole transaction?
Answer #1

Part 1:

If you are going for a short position with the Speculator, what you are doing now is:

  • Deciding the purchase price now for a purchase to be effected after 3 months: P54 per lbs. You will make gain if the price of Oranges per lbs. rises above the purchase price of P54.

Gain = (Market Price - Contracted price) x 36,000 lbs.

  • You will make a loss if the price of Oranges per lbs. falls below the contracted price of P54.

Loss (Contracted price - Market Price) x 36,000 lbs.

Acual situation after 3 months:

  • Price of oranges = P56 per lbs.

1. The price of oranges exceeds the Contracted purchase price of P54, therefore you will make a gain. Therefore our gain will be P72,000.

Gain = (Market Price - Contracted price) x 36,000 lbs.

Gain = (P56 - P54) x 36,000 lbs.= P72,000

2. Loss of Speculator = P72,000

Loss (Contracted price - Market Price) x 36,000 lbs.

Loss (P54 - P56) x 36,000 lbs.= P72,000

Part 2:

The specualtor in order to hedge against the rising prices goes for a short hedge with the hedger at P56. The actual price after 3 months is P57.

1. Gain of Specualtor: P36,000

Gain = (Market Price - Contracted price) x 36,000 lbs.

Gain = (P57 - P56) x 36,000 lbs.= P36,000

2. Loss of the Hedger: P36,000

Loss (Contracted price - Market Price) x 36,000 lbs.

Loss (P56-P57) x 36,000 lbs.= P36,000

3. Overall gain or loss of the speculator: P36,000 Loss

(Loss of P72,000 - Gain of P36,000)